# reference.bib

### v2

@@ Abstract
State-of-the-art large language models (LLMs) are commonly deployed as online services, necessitating users to transmit informative prompts to cloud servers, thus engendering substantial privacy concerns. In response, we present ConfusionPrompt, a novel private LLM inference framework designed to obfuscate the server by: (i) decomposing the prompt into sub-prompts, and (ii) generating pseudo prompts along with the genuine sub-prompts as input to the online LLM. Eventually, the returned responses can be recomposed by the user to obtain the final whole response. Such designs endows our framework with advantages over previous protocols that (i) it can be seamlessly integrated with existing black-box LLMs, and (ii) it achieves significantly better privacy-utility trade-off than existing text perturbation-based methods. We develop a $(\lambda,\mu,\rho)$-privacy model to formulate the requirement for a privacy-preserving group of prompts, and provide a complexity analysis, affirming ConfusionPrompt’s efficiency. Our empirical evaluation reveals that our method offers significantly higher utility compared to local inference methods using open-source models and perturbation-based techniques, while also requiring much less memory than open-source LLMs.
@@ 1 Introduction
Large language models (LLMs) have shown immense potential in various tasks [brown2020language, roziere2023code, huang2023instruct2act]. As a result, they are progressively being incorporated into various real-world applications, including medical consultation [thirunavukarasu2023large] and financial service [wu2023bloomberggpt]. Currently, state-of-the-art LLMs such as ChatGPT from OpenAI [ouyang2022training] are typically provided as online services, since LLM owners deserve the intellectual property of their models and thus they are unwilling to reveal model parameters to the clients [openai2023gpt4]. However, direct usage of such online service could raise significant privacy concerns, since the users’ prompts are likely to hold sensitive information that should remain concealed from the server.
@@
Available solutions to privacy-preserving model inference include encryption [liu2023llms, chen2022x] and perturbation [du2023dp] techniques, which are impractical for private LLM inference. In encryption methods, the server leverages homomorphic encryption (HE) [acar2018survey] and Secure Multiparty Computation (MPC) [cramer2015secure] to conduct private inference on the users’ encrypted query. The practicality of this method is limited by its high computation overheads, especially on LLMs (e.g., GPT-3 has 175B parameters [brown2020language]). The basic idea of a perturbation-based method is to inject a specific level of noise to user’s input before releasing it to the server, which is challenging to strike a nice privacy-utility balance [lyu2020differentially].
@@
Another issue with the above solutions is that they rely on the service provider to modify their framework or release specific model parameters, which could be extremely costly. The application of encryption techniques relies on platform’s investment in hardware and algorithm acceleration. Perturbation-based methods often rely on local differential privacy (LDP) [dwork2006differential] to ensure privacy with formal guarantees. However, recent studies have proposed deploying certain modules on the user side for improved LDP performance, necessitating the sharing of specific model parameters by the server [qu2021natural, mai2023split, malekzadeh2024salted]. These requirements impose additional burdens and proprietary concerns on the service provider, hindering the deployment of these solutions.
@@
To alleviate the above concerns, this paper proposes a novel private LLM inference framework ConfusionPrompt that can be seamlessly integrated with existing online black-box LLMs. The basic idea of ConfusionPrompt is to construct a group of prompts (containing real and fake prompts) with privacy guarantee to confuse the server, making the curious server fail to infer the real user prompt. However, a key challenge arises when a prompt contains multiple private attributes. In such cases, the user needs to consider all combinations of true and fake attributes to effectively confuse an attacker with prior knowledge of certain attributes, rendering the query complexity exponential in the attribute size. To address this issue, we propose a decomposition approach in which the query is divided into sub-questions. This ensures that the private information is distributed across different sub-queries, thereby reducing the complexity to approximately linear in the attribute size (see Figure 4).
@@
Following the above ideas, our proposed ConfusionPrompt consists of four critical steps: (1) The user decomposes the original prompt into several sub-prompts; (2) For each genuine sub-prompt, the user generates a series of pseudo prompts that obfuscate the sensitive attributes in the genuine sub-prompt, until meeting the requirement of privacy guarantee. (3) User queries online services with a group of prompts, receives the corresponding responses, and retrieves the sub-responses for the genuine sub-prompts. (4) User recompose the sub-responses to obtain the final response. Such a framework empowers the user with independent deployment, eliminating the need for additional privacy-preserving actions from the service provider.
@@
Accordingly, we develop a $(\lambda,\mu,\rho)$-privacy model for ConfusionPrompt to formulate the requirements for a privacy-preserving group of prompts. Through complexity analysis, we show that for a prompt $\boldsymbol{p}$ with $U(\boldsymbol{p})$ private attributes and $\mu$ privacy budget, the basic confusion strategy without decomposition requires to generate $\mathcal{O}\left((1/\mu)^{U(\boldsymbol{p})}\right)$ fake prompts. While our decomposition strategy in ConfusionPrompt can reduce the complexity to $\mathcal{O}\left((1/\mu)U(\boldsymbol{p})\right)$ in the ideal decomposition scenario. Overall, based on the privacy model and complexity analysis, we derive the criteria for an ideal decomposer and generator, and accordingly construct a preference dataset to train them using direct preference optimization (DPO) [rafailov2023direct].
Our contributions are as follows:
@@
(1) We are the first to propose a private LLM inference framework using confusion-based strategy. Our ConfusionPrompt framework can be seamlessly implemented by clients with existing online black-box LLMs such as ChatGPT and Claude with improved privacy-utility trade-off.
(2) To safeguard user’s privacy, we define a privacy model to formulate the requirements of a group of prompts.
(3) We leverage a local decomposition module to reduce the complexity of the prompt group under the same privacy protection level.
@@
(4) Experiments show that our ConfusionPrompt can achieve consistently and significantly better utility than local inference method using open-source models and LDP-based methods. Additionally, ConfusionPrompt requires significantly lower memory cost compared to open-source LLMs.
@@ 2 Related Works
Anonymization-based method: Traditional anonymization techniques rely on Named Entity Recognition (NER) to redact Personally Identifiable Information (PII) such as individuals’ names, social security number, and email addresses [ehrmann2023named, lample2016neural]. Recent research on LLM inference involves sanitizing sensitive items in the input and subsequently de-anonymize the LLM’s returned responses [kan2023protecting, chen2023hide]. The downside of this method is that anonymization fall short in concealing other sensitive information, including verbs and non-named entities [anandan2012t].
Encryption-based method: Cryptonets [gilad2016cryptonets] proposed the first neural network inference on encrypted data using homomorphic encryption (HE). They approximated the non-linear function such as Sigmoid and MaxPooling by polynomials. Iron [hao2022iron] designed specialized and efficient protocols for two types of computationally heavy operations in Transformer-based inference: (i) matrix multiplications, and (ii) complex functions including Softmax, GELU activations, and LayerNorm. To achieve further speedup, [liu2023llms] transformed the high-overhead functions into cryptography-friendly approximations, and finetune the model to maintain accuracy.
Perturbation-based method: Perturbation-based method provide privacy guarantee by inject calibrated noise into the input. Existing research utilizing this method predominantly focus on privacy protection in fine-tuning [kerrigan2020differentially, yu2021differentially] and prompt-tuning phases [duan2023flocks, li2023privacypreserving], while few research investigate the privacy-preserving inference paradigm. A major challenge in perturbation-based private inference is to balance the utility and privacy tradeoff. Recent studies have proposed Text2Text [feyisetan2020privacy] and paraphraser-based approaches [utpala2023locally, mattern2022limits] to privatize text with LDP. Split-N-Denoise (SnD) [mai2023split] deployed the token embedding layer at the client side, and introduced a user-side denoising model to correct the purturbed embedding output for downstream tasks.
@@ 3.1 Overall Design
Denote $G_{s}:\mathcal{V}^{*}\rightarrow\mathcal{V}^{*}$ as the large LM deployed as cloud service such as ChatGPT [ouyang2022training, openai2023gpt4], where $\mathcal{V}^{*}$ is the vocabulary space. Instead of sending raw user prompt to the cloud, our ConfusionPrompt introduces a new privacy-preserving inference paradigm, which makes it significantly hard for any third party to infer the real original user prompt even if it happens to access the information sent from the user. Basically, ConfusionPrompt involves the following six steps:
(1) Decomposition of the original prompt. We first introduce a decomposer $G_{d}:\mathcal{V}^{*}\rightarrow\mathcal{V}^{*}$ that aims to decompose the user’s original prompt $p$ to a sequence of genuine sub-prompts $\boldsymbol{p}=[p_{1},p_{2},...,p_{|\boldsymbol{p}|}]$, resulting in fewer private attributes (i.e., private information) in each sub-prompt.
(2) Generation of pseudo prompts. A pseudo prompt generator $G_{f}:\mathcal{V}^{*}\rightarrow\mathcal{V}^{*}$ is introduced to generate a pseudo prompt given a genuine sub-prompt by replacing specified critical information. By generating multiple pseudo prompts for each genuine sub-prompt and mixing them together as a prompt group, we are able to hide the genuine sub-prompts, increasing the difficulty of being inferred. For the simplicity of privacy analysis, we design to enforce the pseudo prompt to be consistent with the genuine prompt in terms of syntactic structure.
(3) Evaluation of privacy level. To ensure that the genuine prompt is safely hidden in the prompt group, we also need to design some criteria (e.g., semantic irrelevance) to evaluate the usability of generated pseudo prompts. Based on these criteria, we will continue sample and filter the pseudo prompts until the prompt group meet the privacy guarantee, which will be introduced in details in 3.2.
(4) Communication with cloud server. Here, the user sends the prompt group to the server for the usage of large LM on the cloud, while the server subsequently sends back the response group to the user. During this process, we expect that it will be extremely hard for the server to infer the original user prompt according to the prompt group, therefore safeguarding user’s privacy.
(5) Retrieval of interested response. Since the user itself is aware of which sub-prompts are genuine, it can seamlessly retrieve the corresponding sub-responses and discard the responses of pseudo prompts.
(6) Recomposition of sub-responses. Here, we introduce a recomposer $G_{r}:\mathcal{V}^{*}\rightarrow\mathcal{V}^{*}$ that maps a sequence of sub-prompt-response pairs to a final response.
Figure 1: Overview of ConfusionPrompt. (1) Decomposition of the original prompt into sub-prompts. (2) Generation of pseudo prompts. (3) Querying online LLMs to get the sub-responses and retrieving the interested sub-responses. (4) Recomposition of sub-responses to obtain the final response.
@@ 3.2 Privacy Model
Here, we define a $(\lambda,\mu,\rho)$-privacy model to formulate the requirements that the group of pseudo prompts should satisfy for privacy protection, which offers guidance for the evaluation of privacy level and training of models (i.e., decomposer, generator, and recomposer) (i) The fabricated attributes in the pseudo prompt should be semantically irrelevant to the original attributes in genuine prompt, so that the genuine user information can be effectively obfuscated [wu2015constructing]. (ii) The genuine prompt should be obfuscated by sufficient pseudo prompts, and the attacker can not easily identify the real prompt from the combination pattern even with background knowledge. (iii) The pseudo prompt should appear genuine rather than obviously fabricated. Thus it’s crucial to maintain the fluency and reasonability of the pseudo prompts.
@@
Definition 3.1 (Private Attributes).
Denote $\mathcal{U}$ as the attribute space, and $\mathcal{P}$ as the prompt space. Given a sub-prompt list $\boldsymbol{p}=\in\mathcal{P}$, the contained private attributes are denoted as $U(\boldsymbol{p})=\{u_{1},u_{2},...,u_{m}\}$ for $m$ attributes.
Remark 3.2.
For conciseness, the prompt in definition 3.1 could be a full prompt $p$, or a sequence of decomposed prompts with the same intention $\boldsymbol{p}=[p_{1},p_{2},...,p_{|\boldsymbol{p}|}]$.
@@
Definition 3.3 (Attribute-attribute Similarity).
@@
Definition 3.4 (Correspondent Attributes).
@@
For example, given the original question ‘What are the responsibilities of software engineers?’ and private attribute ‘software engineer’, the correspondent attribute in the fake question ‘What are the responsibilities of school teachers?’ would be ‘school teacher’. Based on the correspondent attributes and attribute-attribute similarity, we can define the semantic similarity between two prompts below:
Definition 3.6 (Prompt-prompt Similarity).
@@
We proceed to the definition of significance of an attribute, which is tied to the likelihood of a curious server being able to identify the genuine attribute from a group of genuine and pseudo prompts.
Definition 3.7 (Significance of Single Attribute).
Denote =$\mathbb{P}=\{\boldsymbol{p}_{1},\boldsymbol{p}_{2},...,\boldsymbol{p}_{n}\}$ as a group of prompts. The significance of an attribute $u\in U(\boldsymbol{p}_{i})$ related to a group of prompts $\mathbb{P}$ is defined as:
$\mathrm{Sig}(u,\mathbb{P})=\frac{\sum_{j=1}^{n}H(u,\boldsymbol{p}_{i},\boldsymbol{p}_{j})}{n},$ (2)
@@
Definition 3.8 (Significance of Attribute Set).
Suppose $\boldsymbol{p}_{0}$ denotes the genuine prompt. The significance of attribute set $U(\boldsymbol{p}_{i})$ related to a group of prompts =$\mathbb{P}=\{\boldsymbol{p}_{0},\boldsymbol{p}_{1},...,\boldsymbol{p}_{n}\}$ is defined as:
$\mathrm{Sig}(U(\boldsymbol{p}_{0}),\mathbb{P})$ (4)
@@
Based on definition 3.6, 3.8, and 3.9, we can formulate our proposed $(\lambda,\mu,\rho)$-privacy model as:
Definition 3.10 (User Privacy).
Let =$\mathbb{P}=\{\boldsymbol{p}_{0},\boldsymbol{p}_{1},...,\boldsymbol{p}_{n}\}$ be a group of prompts, where $\boldsymbol{p}_{0}$ is the genuine user prompt and the remaining ones are pseudo prompts. If $\mathbb{P}$ satisfies the following requirements, then it is deemed that it can ensure $(\lambda,\mu,\rho)$-privacy of user prompt $\boldsymbol{p}_{0}$ with respect to discriminator $D$:
@@
  • The users’ sensitive attributes should be obfuscated by sufficient pseudo prompts, i.e., $\rm{Sig}(U(\boldsymbol{p}_{0}),\mathbb{P})\leq\mu$.
@@
In Appendix C, we provide the relationship between our privacy model and the upper bound on the success rate of inference attacks, which states that an curious server can do no better than random guessing if the combination of attributes are uniformly distributed.

### v4

# ConfusionPrompt: Practical Private Inference for Online Large Language Models

@@ Abstract
State-of-the-art large language models (LLMs) are typically deployed as online services, requiring users to transmit detailed prompts to cloud servers. This raises significant privacy concerns. In response, we introduce ConfusionPrompt, a novel framework for private LLM inference that protects user privacy by: (i) decomposing the original prompt into smaller sub-prompts, and (ii) generating pseudo-prompts alongside the genuine sub-prompts, which are then sent to the LLM. The server responses are later recomposed by the user to reconstruct the final output. This approach offers key advantages over previous LLM privacy protection methods: (i) it integrates seamlessly with existing black-box LLMs, and (ii) it delivers a significantly improved privacy-utility trade-off compared to existing text perturbation methods. We also develop a $(\lambda,\mu,\rho)$-privacy model to formulate the requirements for a privacy-preserving group of prompts and provide a complexity analysis to justify the role of prompt decomposition. Our empirical evaluation shows that ConfusionPrompt achieves significantly higher utility than local inference methods using open-source models and perturbation-based techniques, while also reducing memory consumption compared to open-source LLMs.
keywords:
Large language model , privacy-preserving computation , private inference , confusion strategy
journal: Knowledge-Based Systems
[…]
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities across a wide range of tasks [1, 2, 3], leading to their growing integration into real-world applications such as medical consultations [4] and financial services [5]. State-of-the-art LLMs, such as OpenAI’s ChatGPT [6], are predominantly offered as online services. This is mainly due to intellectual property considerations, as model owners are reluctant to disclose their proprietary parameters to users [7]. However, direct usage of such LLM online services could raise significant privacy concerns, as users often transmit prompts containing sensitive information that should ideally remain confidential from the server.
@@
Existing solutions to privacy-preserving model inference, such as encryption [8, 9] and perturbation [10] techniques, are often impractical for private LLM inference. In encryption-based methods, the server leverages homomorphic encryption (HE) [11] and Secure Multiparty Computation (MPC) [12] to conduct private inference on the users’ encrypted query. The practicality of this method is limited by its high computation overheads, especially on LLMs (e.g., GPT-3 has 175B parameters [1]). On the other hand, the basic idea of perturbation-based methods is to inject a specific level of noise into the user’s input before releasing it to the server, which is challenging to strike a satisfied privacy-utility balance [13].
@@
Another limitation of the aforementioned solutions is their dependence on the service provider to modify their infrastructure or disclose specific model parameters, which can be prohibitively costly. Encryption-based techniques require significant investment from platform in hardware and algorithmic acceleration to support efficient computation. Perturbation-based methods often rely on local differential privacy (LDP) [14] to ensure privacy with formal guarantees. For improved LDP performance, recent studies have proposed deploying certain modules on the user side, necessitating the sharing of specific model parameters by the server [15, 16, 17]. These requirements impose additional burdens and proprietary concerns on the service provider, hindering the deployment of these solutions.
@@
To alleviate the above concerns, this paper proposes ConfusionPrompt, a novel private LLM inference framework that can be seamlessly integrated with existing online black-box LLMs. The core idea of ConfusionPrompt is to construct a set of prompts, containing real and fake prompts, designed to confuse the server, making the curious server fail to infer the real user prompt. However, a significant challenge arises when a prompt contains multiple sensitive attributes. In such cases, the user needs to consider all combinations of true and fake attributes to effectively confuse an attacker with prior knowledge of certain attributes, rendering the query complexity growing exponentially with the number of attributes. To address this issue, we propose a decomposition approach in which the query is divided into sub-questions. This ensures that the private information is distributed across different sub-queries, thereby reducing the complexity to approximately linear in the nunber of attributes (see Figure 2).
@@
Building on the aforementioned concepts, our proposed ConfusionPrompt consists of four critical steps: (1) The user decomposes the original prompt into several sub-prompts; (2) For each genuine sub-prompt, the user generates a set of pseudo prompts that obfuscate sensitive attributes in the genuine sub-prompt, continuing this process until the desired level of privacy is achieved. (3) The user sends a group of both genuine and pseudo-prompts to the online service, retrieves the responses, and isolates the sub-responses corresponding to the genuine sub-prompts. (4) Finally, user recomposes the sub-responses to obtain the final result. This framework empowers the user with independent deployment, eliminating the need for additional privacy-preserving actions from the service provider.
@@
Accordingly, we develop a $(\lambda,\mu,\rho)$-privacy model for ConfusionPrompt to formulate the requirements for a privacy-preserving group of prompts. Through complexity analysis, we show that for a prompt $\boldsymbol{p}$ with $U(\boldsymbol{p})$ private attributes and a privacy budget $\mu$, the basic confusion strategy without decomposition requires generating $\mathcal{O}\left((1/\mu)^{U(\boldsymbol{p})}\right)$ fake prompts, while our decomposition strategy in ConfusionPrompt can reduce the complexity to $\mathcal{O}\left((1/\mu)U(\boldsymbol{p})\right)$ in the ideal decomposition scenario. Based on the privacy model and complexity analysis, we derive the criteria for an ideal decomposer and generator, and accordingly design a two-stage training strategy.
Our key contributions are as follows:
  • 1.
@@
    We are the first to propose a private LLM inference framework using confusion-based strategy. Our ConfusionPrompt framework can be seamlessly implemented by clients with existing online black-box LLMs, such as ChatGPT and Claude, providing an improved privacy-utility trade-off. In addition, Our framework can be integrated with Trusted Execution Environment (TEE) [18] solutions. Even in the event of a TEE compromise, user privacy remains safeguarded through the confusion strategy.
  • 2.
    To ensure user privacy, we define a privacy model to formulate the requirements of a group of prompts, including both real and pseudo prompts. Furthermore, we introduce a local decomposition module to reduce the complexity of the prompt group under the same privacy parameters.
  • 3.
@@
    Experiments show that ConfusionPrompt can achieve a consistently and significantly better utility than the local inference method using open-source models and LDP-based methods. Moreover, ConfusionPrompt significantly reduces memory overhead compared to open-source LLMs.
@@ 2 Related Works
Anonymization-based method. Traditional anonymization techniques rely on Named Entity Recognition (NER) to redact Personally Identifiable Information (PII) such as individuals’ names, social security number, and email addresses [19, 20]. Recent research on LLM inference involves sanitizing sensitive items in the input and subsequently de-anonymize the LLM’s returned responses [21, 22]. The downside of this method is that anonymization fall short in concealing other sensitive information, including verbs and non-named entities [23].
Encryption-based method. Cryptonets [24] proposed the first neural network inference on encrypted data using homomorphic encryption (HE). They approximated the non-linear function such as Sigmoid and MaxPooling by polynomials. Iron [25] designed specialized and efficient protocols for two types of computationally heavy operations in Transformer-based inference: (i) matrix multiplications, and (ii) complex functions including Softmax, GELU activations, and LayerNorm. To achieve further speedup, [8] transformed the high-overhead functions into cryptography-friendly approximations, and finetuned the model to maintain accuracy.
Perturbation-based method. Perturbation-based methods provide privacy guarantee by inject calibrated noise into the input. Existing studies utilizing this method predominantly focus on privacy protection in fine-tuning [26, 27] and prompt-tuning phases [28, 29], while few studies investigate the privacy-preserving inference paradigm. A major challenge in perturbation-based private inference is to balance the utility and privacy tradeoff. Recent studies have proposed Text2Text [30] and paraphraser-based approaches [31, 32] to privatize text with LDP. Split-N-Denoise (SnD) [16] deployed the token embedding layer at the client side, and introduced a user-side denoising model to correct the purturbed embedding output for downstream tasks.
@@ 3.1 Overall Design
Denote $G_{s}:\mathcal{V}^{*}\rightarrow\mathcal{V}^{*}$ as the LLM deployed as cloud service such as ChatGPT [6, 7], where $\mathcal{V}^{*}$ is the vocabulary space. Instead of sending raw user prompt to the cloud, our ConfusionPrompt introduces a new privacy-preserving inference paradigm, which makes it significantly hard for any third party to infer the real original user prompt even if it happens to access the information sent from the user. Basically, ConfusionPrompt involves the following six steps:
Step 1: decomposition of the original prompt. We first introduce a decomposer $G_{d}:\mathcal{V}^{*}\rightarrow\mathcal{V}^{*}$ that aims to decompose the user’s original prompt $p$ to a sequence of genuine sub-prompts $\boldsymbol{p}=[p_{1},p_{2},...,p_{|\boldsymbol{p}|}]$, resulting in fewer private attributes (i.e., private information) in each sub-prompt.
Step 2: generation of pseudo prompts. A pseudo prompt generator $G_{f}:\mathcal{V}^{*}\rightarrow\mathcal{V}^{*}$ is introduced to generate a pseudo prompt given a genuine sub-prompt by replacing specified critical information. By generating multiple pseudo prompts for each genuine sub-prompt and mixing them together as a prompt group, we are able to hide the genuine sub-prompts, thus increasing the difficulty of being inferred. For the simplicity of privacy analysis, we design to enforce the pseudo prompt to be consistent with the genuine prompt in terms of syntactic structure.
Step 3: evaluation of privacy level. To ensure that the genuine prompt is safely hidden in the prompt group, we also need to design some criteria (e.g., semantic irrelevance) to evaluate the usability of generated pseudo prompts. Based on these criteria, we iteratively sample and filter the pseudo prompts until the prompt group meets the privacy requirement, which will be introduced in details in Section 3.2.
Step 4: communication with cloud server. Here, the user sends the prompt group to the server for the usage of LLM on the cloud, while the server subsequently sends back the response group to the user. During this process, we expect that it will be hard for the server to infer the original user prompt according to the prompt group, therefore safeguarding user’s privacy.
Step 5: retrieval of interested response. Since the user itself is aware of which sub-prompts are genuine, it can seamlessly retrieve the corresponding sub-responses and discard the responses of pseudo prompts.
Step 6: recomposition of sub-responses. Here, we introduce a recomposer $G_{r}:\mathcal{V}^{*}\rightarrow\mathcal{V}^{*}$ that maps a sequence of sub-prompt-response pairs to a final response.
Figure 1: Overview of ConfusionPrompt.
@@ 3.2 Privacy Model
3.2.1 Rationale of Privacy Model
Here, we define a $(\lambda,\mu,\rho)$-privacy model to formulate the requirements that the group of pseudo prompts should satisfy for privacy protection, which offers guidance for the evaluation of privacy level and training of models (i.e., decomposer, generator, and recomposer). To explain the rationale of our privacy model, we follow [33] to quantify the privacy risk of the queries exposed to the server.
Consider a set of prompts denoted as $\boldsymbol{P}=\{\boldsymbol{p}_{1},\boldsymbol{p}_{2},...,\boldsymbol{p}_{n}\}$. For any $\boldsymbol{p}\in\boldsymbol{P}$, let $\pi(\boldsymbol{p})$ be the adversary’s prior probability that $\boldsymbol{p}$ is the genuine prompt provided by the user. Then the adversary’s posterior probability $\pi(\boldsymbol{p}|\boldsymbol{P})$ of the same event is:
$\pi(\boldsymbol{p}|\boldsymbol{P})=\frac{\pi(\boldsymbol{p})}{\sum_{\boldsymbol{p}^{\prime}\in\boldsymbol{P}}\pi(\boldsymbol{p}^{\prime})}.$ (1)
Let $\boldsymbol{p}^{0}$ denote the genuine prompt provided by the user. Then the privacy risk associated with the set of prompts $\boldsymbol{P}$ revealed to the server can be formulated as:
$\mathrm{Risk}(\boldsymbol{P})=\sum_{\boldsymbol{p}^{\prime}\in\boldsymbol{P}}\pi(\boldsymbol{p}^{\prime}|\boldsymbol{P})\cdot\mathrm{Sim}(\boldsymbol{p}^{\prime},\boldsymbol{p}^{0}),$ (2)
where $\mathrm{Sim}$ is some measure of similarity between two prompts.
Equation 2 suggests three possible methods to minimize the privacy risk. First, reducing $\mathrm{Sim}(\boldsymbol{p}^{\prime},\boldsymbol{p}^{0})$ for any $\boldsymbol{p}^{\prime}\neq\boldsymbol{p}^{0}$ leads to a decrease in $\mathrm{Risk}(\boldsymbol{P})$. Therefore, generating deceptive prompts that are semantically distinct from the genuine prompt can effectively lower the privacy risk. Second, $\mathrm{Risk}(\boldsymbol{P})$ is reduced when $\pi(\boldsymbol{p}^{\prime}|\boldsymbol{P})$ is high for prompts $\boldsymbol{p}^{\prime}$ that do not resemble $\boldsymbol{p}^{0}$. This implies that fabricated prompts should be crafted to appear as realistic as possible. Finally, $\mathrm{Risk}(\boldsymbol{P})$ decreases if we increase the number of prompts $\boldsymbol{p}^{\prime}$ with low $\mathrm{Sim}(\boldsymbol{p}^{\prime},\boldsymbol{p}^{0})$. This can be achieved by increasing the number of fake prompts with different semantic meaning.
Based on the observations, we develop the following criteria for the fabricated prompts: (i) The fabricated attributes in the pseudo prompt should be semantically irrelevant to the original attributes in genuine prompt, so that the genuine user information can be effectively obfuscated [34]. (ii) The genuine prompt should be obfuscated by sufficient number of pseudo prompts, and the attacker can not easily identify the real prompt from the combination pattern even with background knowledge. (iii) The pseudo prompt should appear genuine rather than obviously fabricated. Thus it’s crucial to maintain the fluency and reasonability of the pseudo prompts.
3.2.2 Construction of Privacy Model
@@
Definition 1 (Private Attributes).
Denote $\mathcal{U}$ as the attribute space, and $\mathcal{P}$ as the prompt space. Given a sub-prompt list $\boldsymbol{p}\in\mathcal{P}$, its private attributes are denoted as $U(\boldsymbol{p})=\{u_{1},u_{2},...,u_{m}\}$ for $m$ attributes.
Remark 2.
For conciseness, the prompt in Definition 1 could be a full prompt $p$, or a sequence of decomposed prompts with the same intention $\boldsymbol{p}=[p_{1},p_{2},...,p_{|\boldsymbol{p}|}]$.
@@
Definition 3 (Attribute-attribute Similarity).
@@
Definition 4 (Correspondent Attributes).
@@
For example, given the original question ”What are the responsibilities of software engineers?” and private attribute ”software engineer”, the correspondent attribute in the fake question ”What are the responsibilities of school teachers?” would be ”school teacher”. Based on the correspondent attributes and attribute-attribute similarity, we can define the semantic similarity between two prompts below:
Definition 6 (Prompt-prompt Similarity).
@@
We proceed to the definition of significance of an attribute, which is tied to the likelihood of a curious server being able to identify the genuine attribute from a group of genuine and pseudo-prompts.
Definition 7 (Significance of Single Attribute).
Denote $\boldsymbol{P}=\{\boldsymbol{p}_{1},\boldsymbol{p}_{2},...,\boldsymbol{p}_{n}\}$ as a group of prompts. The significance of an attribute $u\in U(\boldsymbol{p}_{i})$ related to a group of prompts $\boldsymbol{P}$ is defined as:
$\mathrm{Sig}(u,\boldsymbol{P})=\frac{\sum_{j=1}^{n}H(u,\boldsymbol{p}_{i},\boldsymbol{p}_{j})}{n},$ (4)
@@
Definition 8 (Significance of Attribute Set).
Suppose $\boldsymbol{p}_{0}$ denotes the genuine prompt. The significance of attribute set $U(\boldsymbol{p}_{i})$ related to a group of prompts $\boldsymbol{P}=\{\boldsymbol{p}_{0},\boldsymbol{p}_{1},...,\boldsymbol{p}_{n}\}$ is defined as:
$\mathrm{Sig}(U(\boldsymbol{p}_{0}),\boldsymbol{P})$ (6)
@@
Based on Definition 6, 8, and 9, we can formulate our proposed $(\lambda,\mu,\rho)$-privacy model as:
Definition 10 (User Privacy).
Let $\boldsymbol{P}=\{\boldsymbol{p}_{0},\boldsymbol{p}_{1},...,\boldsymbol{p}_{n}\}$ be a group of prompts, where $\boldsymbol{p}_{0}$ is the genuine user prompt and the remaining ones are pseudo prompts. If $\boldsymbol{P}$ satisfies the following requirements, then it is deemed that it can ensure $(\lambda,\mu,\rho)$-privacy of user prompt $\boldsymbol{p}_{0}$ with respect to discriminator $D$:
  • 1.
@@
  • 2.
    The users’ sensitive attributes should be obfuscated by sufficient pseudo prompts, i.e., $\rm{Sig}(U(\boldsymbol{p}_{0}),\boldsymbol{P})\leq\mu$.
  • 3.
@@
In Appendix B, we provide the relationship between our privacy model and the upper bound on the success rate of inference attacks, which states that an curious server can do no better than random guessing if the combination of attributes are uniformly distributed.